OpenStudy (anonymous):

The surface area "S" of a right prism is given by S=2B+Ph where "B" is the area of the base, "P" is the perimeter of the base, and "h" is the height of the prism. Solve for "h"?

OpenStudy (anonymous):

S = 2B + Ph S - 2B = Ph (S-2B/P = h

OpenStudy (anonymous):

S=2B+Ph S-2B=Ph (S-2B)/P=h then just put the values and get the answer if you have them.

OpenStudy (anonymous):

(S-2B)/P i mean basically, move all terms to one side and try to get p all by its lonesome

OpenStudy (anonymous):

try to get h* man im not with it